OpenStudy (anonymous):

The resistances of commercially-available discrete resistors are restricted to particular sets. For example, the available values of resistors with 10% tolerance are selections from the E12 set multiplied by a power of ten from 10^0 through 10^5. The E12 set is E12={10, 12, 15, 18, 22, 27, 33, 39, 47, 56, 68, 82}

OpenStudy (anonymous):

Vout = R2/(R2+R1) * Vin, Thus Ratio=Vout /Vin= R2/(R2+R1) With given voltage values here Ratio=1/4 To find Rout , one needs to suppress the independent source, so we short out the voltage source. Therefore Rout=R1||R2 So the requirements of the question are: • Rout min < Rout = R1*R2/(R1+R2) < Rout max (1) • Ratio*(1-ε) < Vout /Vin = R2/(R1+R2) < Ratio*(1+ε) (2) By inspecting inequalities (1) and (2) closely, one can notice if we divide the two, a bound for R1 can be found, so R1 can be selected independent of R2, which is helpful in simplifying the design. However, there is a subtle point in dividing ranges. Note that • If a < x < b and c < y < d • Then a/d < x/y < b/c (Not the simple division of bounds of a/c < x/y < b/d) With this in mind, the bounds for R1 can be found by dividing inequalities of (1) with (2): Rout min / (Ratio*(1+ε)) < R1 < Rout max / (Ratio*(1-ε)) With the values of this problem (Ratio=0.25, ε=0.1, Rout max = 30k and Rout min =10k), R1 bounds are as follows: 10k/0.275 < R1 < 30k/0.225, which results in: 36.4k < R1 < 133.3k With such a bound for R1, there are many options in E12 set that could work (39k, 47k, 56k, 68k, 82k, 100k, 120k are all possible). However for a conservative design, I'll choose a value mid-range, for instance 68k. Therefore R1=68kΩ. With the value of R1 selected, now we can choose R2 such that Vout /Vin ratio satisfies the nominal value: Vout /Vin = R2/(R1+R2) = 1/4 in this case. Therefore 3*R2=R1 and thus R2=22.67k. However we are limited to E12 possible resistor values. Thus I choose the closest value of 22k for R2. However since this is not the exact value derived from Vout /Vin ratio, let's check to see if our design satisfies the 10% tolerance limit on Vout /Vin . With • R1=68 kΩ • R2=22 kΩ Vout /Vin = 0.244 which is well within 10% deviation of 0.25 nominal value of the desired ratio. So our design is confirmed to satisfy all requirements! J In order to find Vout min and Vout max with 10% deviation of resistor values, note that maximizing the numerator and minimizing the denominator, maximizes a ratio. • Vout max = 1.1R2 / (0.9R1+1.1R2) * Vin = 1.1*22k / (0.9*68k + 1.1*22k) * 80 = 0.283 * 80 = 22.67 V • Vout min = 0.9R2 / (1.1R1+0.9R2) * Vin = 0.9*22k / (1.1*68k + 0.9*22k) * 80 = 0.209 * 80 = 16.74 V :))))))

OpenStudy (anonymous):

for different values!!!!!